OpenStudy (anonymous):

Write the following as a ratio in the simplest form: 15 winners out of 120 raffle entries

OpenStudy (anonymous):

50 + 50 + 20 = 120

OpenStudy (anonymous):

better: 15 , 30 , 45 , .....120

OpenStudy (texaschic101):

15/120 = 8/1